# Identifying Weight-Variant Latent Causal Models

### v6

@@ 8 Conclusion
Identifying latent causal representations is known to be generally impossible without certain assumptions. Motivated by recent progress in the identifiability result of nonlinear ICA, we analyse three intrinsic indeterminacies in latent space, e.g., transitivity, permutation indeterminacy and scaling indeterminacy, which provide deep insights for identifying latent causal representations. To address the unidentifiable issue due to transitivity, we explore a new line of research in the identifiability of latent causal representations, that of allowing the causal influences among the latent causal variables to change. We show that latent causal representations, modeled by weight-variant linear Gaussian models, can be identified up to permutation and scaling, with mild assumptions. We then propose a novel method to consistently learn latent causal representations. Experiments on synthetic and fMRI data demonstrate the identifiability and consistency results and the efficacy of the proposed method.
@@ A.1 The proof of Theorem 1
where $\mathbf{I}$ denotes the identity matrix. Due to DAG constraints in a causal system and the assumption of linear models on $\mathbf{z}$, we have 1) the determinant of the Jacobian matrix is equal to 1, e.g., $|\det{\mathbf{J}}_{\mathbf{g_{u}}^{-1}}(\mathbf{z})|=|\det{\mathbf{J}}_{\mathbf{g_{u}}}(\mathbf{u})|=1$, 2) the mapping $\mathbf{g}_{\mathbf{n}}$ is invertible.
Step I: Suppose we have two sets of parameters $\theta=({\mathbf{f},\mathbf{T}_{\mathbf{n}},\boldsymbol{\lambda},\boldsymbol{\beta}},\boldsymbol{\varepsilon})$ and $\hat{\theta}=({\mathbf{\hat{f}},\mathbf{\hat{T}}_{\mathbf{n}},\boldsymbol{\hat{\lambda}},\boldsymbol{\hat{\beta}}})$ such that $p_{({\mathbf{f},\mathbf{T}_{\mathbf{n}},\boldsymbol{\lambda},\boldsymbol{\beta}},\boldsymbol{\varepsilon})}(\mathbf{x}|\mathbf{u})=p_{({\mathbf{\hat{f}},\mathbf{\hat{T}}_{\mathbf{n}},\boldsymbol{\hat{\lambda}},\boldsymbol{\hat{\beta}}})}(\mathbf{x}|\mathbf{u})$ for all pairs $(\mathbf{x},\mathbf{u})$. Due to the assumptions (i) and (ii), by expanding these expressions via the change of variables formula and taking the logarithm we find:
@@
According to lemma 3 in (Khemakhem et al., 2020) and Proof for Theorem 1 IN (Sorrenson et al., 2020) we can show that $\mathbf{A}_{\mathbf{n}}$ has full rank. Since we assume the noise to be Gaussian, Eq. (24) can be re-expressed as:
@@
Then, by the contradiction between 1) for every $n_{i}$, there exists a polynomial with a degree at most 2, and 2) for every $n^{2}_{i}$, there also exists a polynomial with a degree at most 2 (Sorrenson et al., 2020), we can obtain:
@@
where $\mathbf{P}$ denote the permutation matrix with scaling. More details can be found in (Sorrenson et al., 2020).
Step II: Again, suppose we have two sets of parameters $\theta=({\mathbf{f},\mathbf{T}_{\mathbf{z}},\boldsymbol{\lambda},\boldsymbol{\beta}})$ and $\hat{\theta}=({\mathbf{\hat{f}},\mathbf{\hat{T}}_{\mathbf{z}},\boldsymbol{\hat{\lambda}},\boldsymbol{\hat{\beta}}})$ such that $p_{({\mathbf{f},\mathbf{T}_{\mathbf{z}},\boldsymbol{\lambda},\boldsymbol{\beta}})}(\mathbf{x}|\mathbf{u})=p_{({\mathbf{\hat{f}},\mathbf{\hat{T}}_{\mathbf{z}},\boldsymbol{\hat{\lambda}},\boldsymbol{\hat{\beta}}})}(\mathbf{x}|\mathbf{u})$ for all pairs $(\mathbf{x},\mathbf{u})$. Due to the assumptions (i) and (ii), by expanding these expressions via the change of variables formula and taking the logarithm we have:
$\log{|\det\mathbf{J}_{\mathbf{f}^{-1}}(\mathbf{x})|}+\log p_{\mathbf{(\mathbf{T}_{\mathbf{z}},\boldsymbol{\lambda},\boldsymbol{\beta})}}({\mathbf{z}|\mathbf{u}})+\log p_{\boldsymbol{\varepsilon}}(\boldsymbol{\varepsilon})=\log{|\det\mathbf{J}_{\mathbf{\hat{f}}^{-1}}(\mathbf{x})|}+\log p_{\mathbf{(\mathbf{\hat{T}}_{\mathbf{z}},\boldsymbol{\hat{\lambda}},\boldsymbol{\hat{\beta}})}}({\mathbf{\hat{z}}|\mathbf{u}}),$ (27)
@@
where $\mathbf{z}$ denotes $[z_{1},...,z_{i}]$, $\mathbf{z}^{2}$ denotes $[z^{2}_{1},...,z^{2}_{i}]$, $\mathbf{z}_{i\neq j}$ denotes the vector whose elements are ${z}_{i}{z}_{j}$ for all $i\neq j$, $\mathbf{A}_{\mathbf{z}}$ in block matrix can be rewritten as:
@@
${z}_{i}^{2}=$ $\underbrace{(\sum_{j}({A}_{i,j}^{(2)}{\hat{z}^{2}}_{j}))^{2}}_{\mathbf{(a)}}+\underbrace{(\sum_{j}({A}_{i,j}^{(1)}{\hat{z}}_{j}))^{2}}_{\text{(b)}}+\underbrace{({\mathbf{A}}_{i,:}^{(3)}\hat{\mathbf{z}}_{i\neq j})^{2}}_{\text{(c)}}+({c}_{i}^{(1)})^{2}+.....$ (41)
@@
$\left(\begin{array}[]{cccc}s_{1,1}&0&0&...\\ -s_{2,2}\lambda_{1,2}(\mathbf{u})&s_{2,2}&0&...\\ -s_{3,3}\lambda_{1,3}(\mathbf{u})&-s_{3,3}\lambda_{2,3}(\mathbf{u})&s_{3,3}&...\\ .&.&.&...\\ \end{array}\right)$ (54)
$=\left(\begin{array}[]{cccc}A_{1,1}&0&0&...\\ A_{2,1}-A_{1,1}\hat{\lambda}_{1,2}(\mathbf{u})&A_{2,2}&0&...\\ A_{3,1}-\sum_{i=1}^{2}A_{i,1}\hat{\lambda}_{i,3}(\mathbf{u})&A_{3,2}-A_{2,2}\hat{\lambda}_{2,3}(\mathbf{u})&A_{3,3}&...\\ .&.&.&...\\ \end{array}\right).$
@@
Our first assumption is implicitly enforcing Gaussian distribution on the latent noise variables. Note that the assumptions of nonlinear ICA (Khemakhem et al., 2020) on the noise could be broad exponential family distribution, e.g., Gaussian distributions, Laplace, Gamma distribution, and so on. This work considers Gaussian distribution, mainly because it can be straightforwardly implemented for the re-parameterization trick in VAE. 2) it is convenient for simplifying the proof process. It is worthwhile and promising to extend the Gaussian assumption to exponential family for future work.
@@
We allow changes of weight (e.g., causal influences) among latent causal variables across $\mathbf{u}$ to handle the transitivity problem. We argue that this is not a necessary condition for identifiability (Existing methods, including sparse graph structure and temporal information, could be regarded as feasible ways to handle the transitivity.), but it is a sufficient condition to handle the transitivity, and provide a new research line for causal representation learning. In fact, exploring the change of causal influences is not new in observed space (Ghassami et al., 2018; Huang et al., 2020). In addition, identifying latent causal representation by randomly chosen unknown hard intervention Brehmer et al. (2022) can also be regarded as a special change of causal influences. For this viewpoint, we argue that exploring the change of causal influences for identifying latent causal representation is a promising way and not too restricted in reality.
We emphasize that we require the change of weights among latent causal variables across $\mathbf{u}$. But this does not imply that the weights always change across $\mathbf{u}$. That is, some weights are allowed to keep unchanged for some points $\mathbf{u}$, as long as these weights change for other points $\mathbf{u}^{\prime}$. See experimental results in section 7.1.1 for further discussion.
@@ A.4 The Proof of Corollary 3
In this case, the module $p(z_{2})$ is dependent of the module $p(z_{1}|z_{2})$, e.g., both modules depend on the same parameter $\lambda_{1,2}({\mathbf{u}})$, which violates ICM principle. Clearly, the scaling indeterminacy for the recovered latent variables does not affect such dependence. As a result, regardless of the scaling, with the help of ICM, we can fully identify the acyclic causal structure among latent variables $\mathbf{z}$.
@@ A.6.1 Experiments for Figure 7
For experimental results of Figure 7 the number of segments (e.g.,$\mathbf{u}$) $M$ is 10,20,30,40 and 50 respectively. For each segment, the number of latent causal variables is 2 and the sample size is 1000. We consider the following structural causal model
$n_{1}:={\cal N}({\beta}(\mathbf{u})_{1,1},{\beta}(\mathbf{u})_{1,2}),\qquad n_{2}:={\cal N}({\beta}(\mathbf{u})_{2,1},{\beta}(\mathbf{u})_{2,2}),$ (58)
@@
where we sample the mean ${\beta}(\mathbf{u})_{i,1}$ and variance ${\beta}(\mathbf{u})_{i,2}$ from uniform distributions $[-2,2]$ and $[0.01,3]$, respectively. We sample the weights $\lambda_{1,2}(\mathbf{u})$ from uniform distributions $[0.1,2]$. We sample latent variable z according to Eqs. 58 and 59, and then mix them using a 2-layer multi-layer perceptron (MLP) to generate observed $\mathbf{x}$.
@@
For all methods, we used a encoder, e.g. 3-layer fully connected network with 30 hidden nodes and Leaky-ReLU activation functions, and decoder, e.g. 3-layer fully connected network with 30 hidden nodes and Leaky-ReLU activation functions. We also use 3-layer fully connected network with 30 hidden nodes and Leaky-ReLU activation functions. For optimization, we use Adam optimizer with learning rate $31r=1e-$ 3. For hyper-parameters, we set $\beta=4$ for the $\beta$-VAE. For CausalVAE, we set the hyper-parameters as suggested in the paper (Yang et al., 2021).
@@ A.6.2 Experiments for Figures 6 and 8
For experimental results of Figures 6 and 8, the number of segments is $30$ and sample size is 1000, while the number (e.g., dimension) of latent causal variables is 2,3,4,5 respectively. For 2-dimensional latent causal variables, the setting is the same as experiments for Figure 7 (a). For 3-dimensional case, we consider the following structural causal model:
$n_{1}:={\cal N}({\beta}(\mathbf{u})_{1,1},{\beta}(\mathbf{u})_{1,2}),\quad n_{2}:={\cal N}({\beta}(\mathbf{u})_{2,1},{\beta}(\mathbf{u})_{2,2}),\quad n_{3}:={\cal N}({\beta}(\mathbf{u})_{3,1},{\beta}(\mathbf{u})_{3,2}),$ (60)
@@
$n_{1}:={\cal N}({\beta}(\mathbf{u})_{1,1},{\beta}(\mathbf{u})_{1,2}),\qquad n_{2}:={\cal N}({\beta}(\mathbf{u})_{2,1},{\beta}(\mathbf{u})_{2,2}),$ (62)
$n_{3}:={\cal N}({\beta}(\mathbf{u})_{3,1},{\beta}(\mathbf{u})_{3,2}),\qquad n_{4}:={\cal N}({\beta}(\mathbf{u})_{4,1},{\beta}(\mathbf{u})_{4,2}),$ (63)
@@
$n_{1}:={\cal N}({\beta}(\mathbf{u})_{1,1},{\beta}(\mathbf{u})_{1,2}),\quad n_{2}:={\cal N}({\beta}(\mathbf{u})_{2,1},{\beta}(\mathbf{u})_{2,2}),$ (66)
$n_{3}:={\cal N}({\beta}(\mathbf{u})_{3,1},{\beta}(\mathbf{u})_{3,2}),\quad n_{4}:={\cal N}({\beta}(\mathbf{u})_{4,1},{\beta}(\mathbf{u})_{4,2}),\quad n_{5}:={\cal N}({\beta}(\mathbf{u})_{5,1},{\beta}(\mathbf{u})_{5,2}),$ (67)
@@
Again, for all these Eqs 60-69, we sample the mean ${\beta}(\mathbf{u})_{i,1}$ and variance ${\beta}(\mathbf{u})_{i,2}$ from uniform distributions $[-2,2]$ and $[0.01,3]$, respectively. We sample the weights $\lambda_{i,j}(\mathbf{u})$ from uniform distributions $[0.1,2]$. We sample latent variable z according to these equations, and then mix them using a 2-layer multi-layer perceptron (MLP) to generate observed $\mathbf{x}$.
@@ A.6.3 Experiments for Table 1
For experimental results of Table 1, the number of segments is $30$ and the sample size is 1000, and the number (e.g., dimension) of latent causal variables is 3. For the case of the matching assumptions, we consider the structural causal model as mentioned in Eqs. (60)-(61). For the cases of uniform, Laplace and Gamma noise, we replace the original Gaussian noise by using uniform noise, Laplace and Gamma noise, respectively, while the remaining parts are the same as Eqs. (60)-(61). For the case of unchanged weights, we replace the weights in Eqs. (60)-(61) by unchanged weights across $\mathbf{u}$, which are sampled from an uniform distribution $[0.1,2]$. For the case of violated assumption (v), we replace the weights in Eqs. (60)-(61) by $\lambda_{i,j}(\mathbf{u})+b$, where $b$ is a constant from an uniform distribution $[0.1,2]$. The network architecture, optimization, and hyper-parameters are the same as used for experiments of Figure 7 (a).

### v7

@@ 8 Conclusion
Identifying latent causal variables from observed data is generally impossible without additional assumptions. Motivated by recent advances in the identifiability of nonlinear ICA, we investigate three fundamental sources of indeterminacy in the latent space, i.e., transitivity, permutation, and scaling, which offer key insights into the challenges and opportunities in identifying latent causal variables. Among these, we show transitivity as the primary obstacle to identifiability and propose a new research direction that leverages changes in causal influences between latent variables to overcome it. Under weight-variant linear Gaussian models and mild conditions inspired by nonlinear ICA, we prove that latent causal variables can be identified up to permutation and scaling. Importantly, we show that partial identifiability remains possible even when only a subset of the causal influences varies, making our results more applicable in practice. Building on these theoretical insights, we introduce Structural caUsAl Variational autoEncoder (SuaVE), a novel method for jointly learning latent causal variables and their underlying causal graph. Empirical results on both synthetic and fMRI data validate our identifiability theory and demonstrate the practical effectiveness of the proposed method.
@@
This project was partially funded by the Responsible AI Research Centre (Yuhang Liu, Zhen Zhang, Anton van den Hengel, and Javen Qinfeng Shi). Dong Gong was partially supported by the Australian Government through the Australian Research Council Discovery Early Career Researcher Award (DECRA)(DE230101591). Mingming Gong was partially supported by the Australian Government through the Australian Research Council Discovery Projects (DP240102088). The authors also thank the anonymous reviewers for their constructive feedback.
@@ A.1 The proof of Theorem 1
where $\mathbf{I}$ denotes the identity matrix. Due to DAG constraints in a causal system and the assumption of linear models on $\mathbf{z}$, we have 1) the determinant of the Jacobian matrix is equal to 1, i.e., $|\det{\mathbf{J}}_{\mathbf{g_{u}}^{-1}}(\mathbf{z})|=|\det{\mathbf{J}}_{\mathbf{g_{u}}}(\mathbf{u})|=1$, 2) the mapping $\mathbf{g}_{\mathbf{n}}$ is invertible.
[…]
@@
According to lemma 3 in Khemakhem et al. (2020) and Proof for Theorem 1 in Sorrenson et al. (2020) we can show that $\mathbf{A}_{\mathbf{n}}$ has full rank. Since we assume the noise to be Gaussian, Eq. (24) can be re-expressed as:
@@
Then, by the contradiction between 1) for every $n_{i}$, there exists a polynomial with a degree at most 2, and 2) for every $n^{2}_{i}$, there also exists a polynomial with a degree at most 2 (Sorrenson et al., 2020), we can obtain: $n_{i}=A_{i,j}\hat{n}_{j}+c_{i}$. For simplicity, in the following we neglect the noise term $\boldsymbol{\varepsilon}$ in Eq (3). As a result, we can express the result in vector form as:
@@
where $\mathbf{P}$ denote the permutation matrix with scaling. More details can be found in Sorrenson et al. (2020). Note that this simplification is for convenience only, the identifiability result still holds even when the noise term is included.
[…]
$\log{|\det\mathbf{J}_{\mathbf{f}^{-1}}(\mathbf{x})|}+\log p_{\mathbf{(\mathbf{T}_{\mathbf{z}},\boldsymbol{\lambda},\boldsymbol{\beta})}}({\mathbf{z}|\mathbf{u}})+\log p_{\boldsymbol{\varepsilon}}(\boldsymbol{\varepsilon})=\log{|\det\mathbf{J}_{\mathbf{\hat{f}}^{-1}}(\mathbf{x})|}+\log p_{{(\mathbf{\hat{T}}_{\mathbf{z}},\boldsymbol{\hat{\lambda}},\boldsymbol{\hat{\beta}})}}({\mathbf{\hat{z}}|\mathbf{u}}),$ (27)
@@
where $\mathbf{z}$ denotes $z_{1},...,z_{i}]$, $\mathbf{z}^{2}$ denotes $z^{2}_{1},...,z^{2}_{i}]$, $\mathbf{z}_{i\neq j}$ denotes the vector whose elements are ${z}_{i}{z}_{j}$ for all $i\neq j$, $\mathbf{A}_{\mathbf{z}}$ in block matrix can be rewritten as:
@@
${z}_{i}^{2}=$ $\underbrace{(\sum_{j}({A}_{i,j}^{(2)}{\hat{z}^{2}}_{j}))^{2}}_{\mathbf{(a)}}+\underbrace{(\sum_{j}({A}_{i,j}^{(1)}{\hat{z}}_{j}))^{2}}_{\text{(b)}}+\underbrace{({\mathbf{A}}_{i,:}^{(3)}\hat{\mathbf{z}}_{i\neq j})^{2}}_{\text{(c)}}+({c}_{i}^{(1)})^{2}+..$. (41)
@@
[…]
@@
One of our model assumption is enforcing Gaussian distribution on the latent noise variables. Note that the assumptions of nonlinear ICA (Khemakhem et al., 2020) on the noise could be broad exponential family distribution, e.g., Gaussian distributions, Laplace, Gamma distribution, and so on. This work considers Gaussian distribution, mainly because it can be straightforwardly implemented for the re-parameterization trick in VAE. 2) It is convenient for simplifying the proof process. It is worthwhile and promising to extend the Gaussian assumption to exponential family for future work.
@@
We allow changes of weights (i.e., causal influences) among latent causal variables across $\mathbf{u}$ to handle the transitivity problem. We argue that this may be not a necessary condition for identifiability (Existing methods, including sparse graph structure and temporal information, could be regarded as feasible ways to handle the transitivity.), but it is a sufficient condition to handle the transitivity, and provide a new research line for causal representation learning. In fact, exploring the change of causal influences is not new in observed space (Ghassami et al., 2018; Huang et al., 2020). In addition, identifying latent causal variables by randomly chosen unknown hard intervention can also be regarded as a special change of causal influences (Brehmer et al., 2022). From this viewpoint, we argue that exploring the change of causal influences for identifying latent causal representation may be a promising way and not too restricted in reality.
We emphasize that changes of weights among latent variables across different values of $\mathbf{u}$ are essential for achieving full identifiability. However, this does not mean that all weights must vary continuously or at every point in $\mathbf{u}$. Some weights may remain unchanged at certain values of $\mathbf{u}$, as long as they exhibit variation at other values of $\mathbf{u}$. This flexibility is both theoretically sufficient and practically relevant. See Section 7.1.1 for experimental evidence and further discussion on this point.
@@ A.4 The Proof of Corollary 3
In this case, the module $p(z_{2})$ is dependent of the module $p(z_{1}|z_{2})$, i.e., both modules depend on the same parameter $\lambda_{1,2}({\mathbf{u}})$, which violates ICM principle. Clearly, the scaling indeterminacy for the recovered latent variables does not affect such dependence. As a result, regardless of the scaling, with the help of ICM, we can fully identify the acyclic causal structure among latent variables $\mathbf{z}$.
@@ A.6.1 Experiments for Figure 7
For experimental results of Figure 7 the number of segments (i.e., $\mathbf{u}$) $M$ is 10, 20, 30, 40 and 50 respectively. For each segment, the number of latent causal variables is 2 and the sample size is 1000. We consider the following structural causal model
$n_{1}:={\cal N}({\beta}_{1,1}(\mathbf{u}),{\beta}_{1,2}(\mathbf{u})),\qquad n_{2}:={\cal N}({\beta}_{2,1}(\mathbf{u}),{\beta}_{2,2}(\mathbf{u})),$ (58)
@@
where we sample the mean ${\beta}_{i,1}(\mathbf{u})$ and variance ${\beta}_{i,2}(\mathbf{u})$ from uniform distributions [2,2][-2,2] and $[0.01,3]$, respectively. We sample the weights $\lambda_{1,2}(\mathbf{u})$ from uniform distributions $[0.1,2]$. We sample latent variable z according to Eqs. 58 and 59, and then mix them using a 2-layer multi-layer perceptron (MLP) to generate observed $\mathbf{x}$.
@@
For all methods, we used a encoder, i.e., 3-layer fully connected network with 30 hidden nodes and Leaky-ReLU activation functions, and decoder, i.e., 3-layer fully connected network with 30 hidden nodes and Leaky-ReLU activation functions. We also use 3-layer fully connected network with 30 hidden nodes and Leaky-ReLU activation functions. For optimization, we use Adam optimizer with learning rate 1e31e-3. For hyper-parameters, we set $\beta=4$ for the $\beta$-VAE. For CausalVAE, we set the hyper-parameters as suggested in Yang et al. (2021).
@@ A.6.2 Experiments for Figures 6 and 8
For experimental results of Figures 6 and 8, the number of segments is $30$ and sample size is 1000, while the number of latent causal variables is 2,3,4,5 respectively. For 2-dimensional latent causal variables, the setting is the same as experiments for Figure 7 (a). For 3-dimensional case, we consider the following structural causal model:
$n_{1}$ $:={\cal N}({\beta}_{1,1}(\mathbf{u}),{\beta}_{1,2}(\mathbf{u})),\quad n_{2}:={\cal N}({\beta}_{2,1}(\mathbf{u}),{\beta}_{2,2}(\mathbf{u})),\quad n_{3}:={\cal N}({\beta}_{3,1}(\mathbf{u}),{\beta}_{3,2}(\mathbf{u})),$ (60)
@@
$n_{1}:={\cal N}({\beta}_{1,1}(\mathbf{u}),{\beta}_{1,2}(\mathbf{u})),\qquad n_{2}:={\cal N}({\beta}_{2,1}(\mathbf{u}),{\beta}_{2,2}(\mathbf{u})),$ (62)
$n_{3}:={\cal N}({\beta}_{3,1}(\mathbf{u}),{\beta}_{3,2}(\mathbf{u})),\qquad n_{4}:={\cal N}({\beta}_{4,1}(\mathbf{u}),{\beta}_{4,2}(\mathbf{u})),$ (63)
@@
$n_{1}:={\cal N}({\beta}_{1,1}(\mathbf{u}),{\beta}_{1,2}(\mathbf{u})),\quad n_{2}:={\cal N}({\beta}_{2,1}(\mathbf{u}),{\beta}_{2,2}(\mathbf{u})),$ (66)
$n_{3}:={\cal N}({\beta}_{3,1}(\mathbf{u}),{\beta}_{3,2}(\mathbf{u})),\quad n_{4}:={\cal N}({\beta}_{4,1}(\mathbf{u}),{\beta}_{4,2}(\mathbf{u})),\quad n_{5}:={\cal N}({\beta}_{5,1}(\mathbf{u}),{\beta}_{5,2}(\mathbf{u})),$ (67)
@@
Again, for all these Eqs 60-69, we sample the mean ${\beta}(\mathbf{u})_{i,1}$ and variance ${\beta}(\mathbf{u})_{i,2}$ from uniform distributions [2,2][-2,2] and $[0.01,3]$, respectively. We sample the weights $\lambda_{i,j}(\mathbf{u})$ from uniform distributions $[0.1,2]$. We sample latent variable $\mathbf{z}$ according to these equations, and then mix them using a 2-layer MLP to generate observed $\mathbf{x}$.
@@ A.6.3 Experiments for Table 1
For experimental results of Table 1, the number of segments is $30$ and the sample size is 1000, and the number of latent causal variables is 3. For the case of the matching assumptions, we consider the structural causal model as mentioned in Eqs. (60)-(61). For the cases of uniform, Laplace and Gamma noise, we replace the original Gaussian noise by using uniform noise, Laplace and Gamma noise, respectively, while the remaining parts are the same as Eqs. (60)-(61). The network architecture, optimization, and hyper-parameters are the same as used for experiments of Figure 7 (a).
A.6.4 Implementation of Enforcing Assumption (v)
To enforce assumption (v), which requires $\lambda_{i,j}(\mathbf{u}=\mathbf{0})=0$ in the prior model and $\lambda^{\prime}_{i,j}(\mathbf{u}=\mathbf{0})=0$ in the posterior model, we augment the training process with synthetic pairs that regularize model behavior at $\mathbf{u}=\mathbf{0}$, in the absence of observed data at this point. We construct synthetic pairs of the form: $\text{(input: }\mathbf{u}=\mathbf{0},\quad\text{target: }0\text{)}$, where the target value “0” is not associated with the observed variable $\mathbf{x}$, but instead reflects the expected output of a specific model component evaluated at $\mathbf{u}=\mathbf{0}$. Specifically, for the prior model Eq. (10), this corresponds to enforcing $\lambda_{i,j}(\mathbf{0})=0$. For the inference model Eq. (11), it enforces $\lambda^{\prime}_{i,j}(\mathbf{0})=0$. These synthetic pairs are therefore not part of the data distribution $(\mathbf{u},\mathbf{x})$, but rather encode theoretical constraints derived from assumption (v). We integrate the synthetic pairs into the training objective alongside observed data $(\mathbf{u},\mathbf{x})$. For each pair, we apply an $L_{1}$ penalty on the model output at $\mathbf{u}=\mathbf{0}$: $\lVert\lambda_{i,j}(\mathbf{0})\rVert_{1}$ for the prior model, $\lVert\lambda^{\prime}_{i,j}(\mathbf{0})\rVert_{1}$ for the inference model. These terms are added to the ELBO loss as shown in Eq. (13), enabling the model to conform to assumption (v).